2 8. (10 points) Find the general solution to the homogeneous system of DE: | 3 x = Ax where A = -2 1 - 1 -2 -4.
Answer #1

x= AX 3 2 6 where AS -2 1-2 -2 -4 This is dystem of a differential equation dimultaneous of Ist order . A Eigenvalues of +30+Now, eigenvector Corresponding to 1= 2 [A-21V=0 1 2 6 0 -2 103) :9-10:58:8 - 6 -2 4 + 2 V₂ + 6 V₂ = 0 3V2 + 1013 let = 3,3 V2dit W x= G e* + Gets tv + Getst 1,2**57*1*:) + sep) 1 x 1) t -3t X- gt Ce tre + Ge at -10ce 2Get -cet + 3 Geet Get is the a